OpenStudy (anonymous):

your parents are considering renting yoh an apartment instead of paying room and board at your college. the month to month contract is $539.00 month and the annual contact is $467.00 month. if you break the annual contact there is a 2 month penalty

OpenStudy (anonymous):

So what are you asking?

OpenStudy (anonymous):

if you in to an annual contract but decide to leave after 5 months how much do your parents lose by not doing the month to month

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

So is the cost of each month 539.00 dollars in the month to month contract or is it the whole year. And the annual contract which is 467.00 dollars per month or each year?

OpenStudy (anonymous):

month is 539

OpenStudy (anonymous):

Kay.

OpenStudy (anonymous):

So the parents are doing the annual contract, which would be 467.00 dollars per month, correct?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Okay. Sorry about the time the computer went crazy. So if it's 467.00 per month in an annual contract. The whole year would be 5604.00 dollars. However, 5 months times 467.00 dollar is 2335 dollars.

OpenStudy (anonymous):

The parents would lose 2,335.00 dollars after leaving 5 months. Hope this helps!

OpenStudy (texaschic101):

you forgot the 2 month penalty

OpenStudy (anonymous):

So are we doing the month to month contract or annual contract?

OpenStudy (texaschic101):

month to month is 539 x 12 annual is 467 x 7 (the 5 months stayed + the 2 month penalty)

OpenStudy (anonymous):

Oh! So we're doing annual contract still? Oh! Sorry. :0

OpenStudy (anonymous):

Okay, so the whole amount of money would be 3269.00 total lost?

OpenStudy (anonymous):

so the whole amount of money would be 3269.00 total lost? plzz give him a medal he deserves it!!

OpenStudy (johnweldon1993):

Sorry for a late round up here Month to month \(\large $539 \times 5\text{months}\) Annual \(\large $467 \times 5\text{months} +$467 \times 2 \text{ for the 2 month penalty}\) So the total cost for the month to month would have been $2695.00 And the total cost for the Annual contract would have been $3269.00 So in order to see how much they lost by choosing the annual over the monthly, we subtract the 2 and see \(\large 3269 - 2695 = $574\) was how much more they had to overpay *aka "lose"* by choosing this option

OpenStudy (anonymous):

Great job man! No problem! Every bit counts. ;)

OpenStudy (texaschic101):

I messed up...I was doing 539 for the whole year instead of for 5 months...oops